OpenStudy (anonymous):

If m angle 4 = 2y + 32 and m angle 5 = 3y - 3, find y. MEDAL!!!

OpenStudy (dtan5457):

There must be a diagram to this..

OpenStudy (anonymous):

oh yeah sorry one sec

OpenStudy (dtan5457):

Anytime now..lol

OpenStudy (anonymous):

OS isn't letting me attach any files, but I do know the angles are alternate interior

OpenStudy (dtan5457):

That means they are equal to each other.

OpenStudy (anonymous):

I know that, but I keep confusing myself when I try to figure it out

OpenStudy (dtan5457):

2y+28 3y-8 2y+28=3y-8 solve for y

OpenStudy (anonymous):

thanks for restating my original question

OpenStudy (dtan5457):

i set them together after isolating each term to = 0.

OpenStudy (anonymous):

6.4

OpenStudy (anonymous):

or 4 /2/5

OpenStudy (dtan5457):

Actually, I made have made a mistake 4 = 2y + 32 is the 4 even part of the equation? or is that just a label for the angle?

OpenStudy (dtan5457):

2y+32=3y-3 -2y -2y 32=y-3 y=35

OpenStudy (anonymous):

4 is just the angle, so is 5

OpenStudy (anonymous):

thats what i tried saying uhh

OpenStudy (anonymous):

i was gonna finsh

OpenStudy (dtan5457):

@amysparkly12 well i didn't see how those numbers can help solve the problem. explain?

OpenStudy (anonymous):

Thanks so much! it works out:) @dtan5457

OpenStudy (dtan5457):

np

OpenStudy (anonymous):

@amysparkly12 no, sorry but it wasn't even remotely correct

OpenStudy (anonymous):

actully yes,i use this site its a national most popual math calutator its called mathway

OpenStudy (dtan5457):

this question is fairly simple and doesn't require a calculator. what did you even type in?

OpenStudy (anonymous):

I've also used that website multiple times, but the answer you gave me was not correct.

OpenStudy (dtan5457):

@Danielle_Ello anymore questions? or i'm gonna go :O

OpenStudy (anonymous):

okay no its not a caluator it has solveing and calulator in it to

OpenStudy (anonymous):

thats why its really good

OpenStudy (dtan5457):

Why is the answer wrong though>

OpenStudy (anonymous):

uhh whatever bye,your wrong beacuse i have live math professors on there,that help me with math,and english

OpenStudy (anonymous):

@dtan5457 yeah, if you're still there. Just gimme a second to type them out quick

OpenStudy (dtan5457):

I couldn't care less, lol. The point is, your answers for this question are wrong. Plain and done.

OpenStudy (anonymous):

ha ha ha okay

OpenStudy (dtan5457):

Ok, now go somewhere and asks your professors how to plug in questions on a online solver correctly.

OpenStudy (anonymous):

Find the distance between each pair of points: 7. F(3,6), G(7,-4) 8. X(-2,5), Y(1,11) 9. R(8,0), S(-9,6) 10. A(14,-3), B(9,-9)

OpenStudy (anonymous):

burrrrnn lol anyway I need help with those questions ^ por favor

OpenStudy (dtan5457):

distance formula every thing

OpenStudy (dtan5457):

I'll do the first pair

OpenStudy (anonymous):

Ok, thanks

OpenStudy (anonymous):

and show the steps so I can work out the other ones please

OpenStudy (dtan5457):

F(3,6), G(7,-4) \[\sqrt{(7-3)^2+(-4-6)^2}=\sqrt{(4)^2+(-10)^2}=\sqrt{116}=2\sqrt{29}\]

OpenStudy (dtan5457):

I'm assuming they want simplified radicals or else the answer for that would be 10.7703296143

OpenStudy (anonymous):

okay, so it all makes sense to me except for the end part (square root of 29 with the 2)

OpenStudy (anonymous):

Could I just leave it at 116?

OpenStudy (anonymous):

Couldn't*

OpenStudy (anonymous):

@dtan5457

OpenStudy (dtan5457):

you could, but it's just more simplified at 2 radical 29. how that works is you square the outside number 2^2=4, then multiply by inside radical 116.

OpenStudy (anonymous):

oh, okay I guess that makes sense, but I'm pretty sure i can just leave it as it is (116) also, I was wondering if you come back on later (before tomorrow) can you check my answers for the other 3 problems?? 1. 116 (the one you did) 2. 45 3. 325 4. 61 and of coarse they are all in square root form^

OpenStudy (anonymous):

@dtan5457

OpenStudy (dtan5457):

all correct @Danielle_Ello